OpenStudy (yisrael012332):

A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (yisrael012332):

@SithsAndGiggles

OpenStudy (yisrael012332):

@wolf1728

OpenStudy (yisrael012332):

@JoannaBlackwelder

OpenStudy (yisrael012332):

@Elsa213

OpenStudy (yisrael012332):

@eninone

OpenStudy (yisrael012332):

@EllenJaz17

OpenStudy (yisrael012332):

@princeharryyy

OpenStudy (yisrael012332):

@peachpi

OpenStudy (yisrael012332):

help please....

OpenStudy (princeharryyy):

I can't do the steps as if now, that will take time and I've got a lot of work to do. Sorry about that.

OpenStudy (yisrael012332):

ok

OpenStudy (yisrael012332):

anyone else im so lost

OpenStudy (owlcoffee):

To create a new system of equation with the conditions given, let's take for example the first equation: \[-3x+7y=-16\] And let's forget about the "=" sign and center on the following expression: \[-3x+7y+16\] So I will add to it the second equation, multiplied by any number... let's say... 10... \[(-3x+7y+16)+(10)(-9x+5y-16)\] So, doing some simple algebra: \[(-3x+7y+16)-90x+50y-160\] Ending up with: \[-93x+57y-144\] And for the second part you'll have to solve the new system of equation formed by the new equation: \[-93x+57y=144\] \[-9x+5y=16\]

OpenStudy (yisrael012332):

ty so much